Method of Substitution
7. There are 35 people in a room. There are seven more males than females in the room. How many males are there? How many females?
can you please help me to solve this problem thanks?

Answer by edjones(8007)

Let m=number of males & f=number of females.
m=f+7
m+f=35
f+7+f=35
2f+7=35
2f=28
f=14
m=21
.
Ed